jabez177 (jabez177):

What is the percent decrease from 60 to 39?

jabez177 (jabez177):

40%?

OpenStudy (jackky):

65 i think

OpenStudy (dan815):

60 - x% of 60= 39

OpenStudy (agl202):

U all wrong! The percent decrease we are going to use is "x" If something decreases by 30%, then you are actually multiplying the number by 70%, which is 100% - 30% = 70%. Therefore: 60 (1 - x) = 39 60 - 60x = 39 -60x = -21 x = 0.35 or 35% The percent decreased 35% from 60 to 39

OpenStudy (dan815):

60-(x/100)*60=39 60(1-x/100)=39 solve for x

OpenStudy (jackky):

jk ik that

jabez177 (jabez177):

Oh...

OpenStudy (jackky):

yea srry